M is the midpoint of segment AB. Find the coordinates of B if the coordinates of A are (-2,5) and the coordinates of M are (8,-12).
Answer by lwsshak3(11628)

M is the midpoint of segment AB. Find the coordinates of B if the coordinates of A are (-2,5) and the coordinates of M are (8,-12).
**
midpoint formula: ((x1+x2)/2, (y1+y2)/2)
(8,-12)=((x1-2)/2,(y1+5)/2)
..
(x1-2)/2=8
x1-2=16
x1=18
..
(y1+5)/2=-12
y1+5=-24
y1=-29
...
coordinates of B: (18,-29)
Check:
((x1-2)/2,(y1+5)/2)
=((18-2)/2,(-29+5)/2)
=(16/2,-24/2)
=(8,-12)